Mathematics 122 – FALL 96 CALCULUS II

Section 42076 – MTWRF 2:30 –3:20pm – Room: 301 SNOW

Instructor: Prof. Rodolfo H. Torres.

Office: 623 Snow – TE : 864–7310. E–mail: torresmath.ukans.edu Tentative Office Hours: M 12:30–1:30pm,

W F: 10:30–11:30am, or by appointment.

Textbook: Calculus with Analytic Geometry (4th Edition)

by Edwards and Penney. (Required)

Calculator: TI-85 Graphing Calculator. (Required)

Prerequisites: Math 121 or Math 116.

Enrollment Last day of add/drop: September 4th. Changes: 151 STRONG Hall (All changes: drops, adds).

Withdrawals: • Aug 21– Sep 25 . The course is deleted from your record. • Sep 26 – Nov 1. A grade of W goes on your record.

• Nov 2 – Dec 7. Requires a petition to the College of LAS.

Timetable

Exams  Date
In class Exam 1 Thursday, September 19.
Midterm Exam Thursday, October 17, 5:45–7:45 pm.
In class Exam 2 Thursday, November 9.
Gateway Test Before November 27.
Final Exam  Wednesday, December 18, 7:00–10:00 pm.

There will be no make-up exams or quizzes.

Grading

Your grade in the course will be determined by the number of points that you accumulate. The points are distributed in the following way:

Homework, Quizzes and Projects 200 points In-class Exam 1 and 2 200 points Midterm Exam 200 points Gateway Test 100 points Final Exam 300 points

Total 1000 points

Homeworks

Homework problems will be assigned every class and they will be collected every Monday and Thursday at the beginning of the class. Homework is a major part of the learning process in Mathematics. It is essential that students work on the problems on a regular basis. Students are encouraged to work and discuss the problems in groups, but they should turn in individual solutions.

No late homework assignment will be accepted.

Classes

You are expected to bring your textbook and the calculator every class. You should read the book and attempt to solve some of the problems in preparation for each class.

Gateway Test

This is a test of basic skills in techniques of integration . You will be given a list of problems about integration and the gateway test will consist of 10 problems from this list. You must accurately complete at least 9 of these problems. If you are successful, you will receive 100 points; no partial credit is given. There will be opportunities for repeated attempts of the gateway test.

Topics Covered

After a short review about inverse trigonometric functions and integration techniques, the following chapters and sections of the book will be covered in the order indicated: 11.2–11.4, 10.2, 10.3, 12.1–12.4, 13.4, 13.1–13.7, 14.1–14.10, 15.1–15.4, 15.6, 15.7, 16.1–16.4. If time permits some additional topics will be covered.